Can someone please help me with the following question:
The current I in electrical conductor varies inversely as the resistance R of the conductor. Suppose I is 0.2 amperes when the resistance is 200 ohms. Find the current when the resistance is 250 ohms.

Found 2 solutions by Alan3354, Mathtut:
Answer by Alan3354(69443)

Can someone please help me with the following question:
The current I in electrical conductor varies inversely as the resistance R of the conductor. Suppose I is 0.2 amperes when the resistance is 200 ohms. Find the current when the resistance is 250 ohms.
---------------------------
Varying inversely means as R increases, I decreases. Since it's a linear relationship (it wasn't stated otherwise), I*R = a constant. In actuality, it's voltage, V = I*R.
So, V = 0.2*200 = 40 volts.
For R = 250, V = 40 = I*250
I = 40/250 = 0.16 amperes.

Answer by Mathtut(3670)

when you have and inverse relation as opposed to a direct proportion it means that as one thing increases the other decreases or vise versa. so in a direct relationship y directly related to x for instance you have a
relationship of y/x=k but in an inverse relationship y inversely related to x the y*x=k so in our problem I*R=k in this case
(.2)200=40 now we are asked to find I when R=250 so I(250) must equal our constant number which is 40 so I(250)=40 divide by 250 and you get .16
an other way of doing this is to put the ratio of I/I2=R2/R
so (.2)/I=250/200 I=.16